OpenStudy (anonymous):

What is the surface area of a triangular pyramid if each face has an area of 120 yd2? A. 30 yd2 B. 360 yd2 C. 480 yd2 D. 600 yd2

OpenStudy (unklerhaukus):

how many faces does a triangular pyramid have?

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

LACEY! :D I think it's c. 480

OpenStudy (anonymous):

I got 480 here's how The answer depends as to whether you include the base. The base has three sides. Therefore there are three planes rising up from the base to the centre point. Without the base, there are three planes, which each have a area of 120 square yards 3 times 120 = 360 square yards. With the base, there are four planes, which each have a area of 120 square yards 4 times 120 = 480 square yards.

OpenStudy (anonymous):

Copy cat .-. jkjkjk XD

OpenStudy (anonymous):

I am not a copy cat because you just gave the answer, I, on the other hand, explained.

OpenStudy (anonymous):

S.A. = L.A. + B. = 2bl + b^2 = 2(450)(367) + 450^2 = 330,300 + 202,500 = 532,800. Lateral Area: L.A. = 2bl = 2(450)(367) = 330,300 Hopefully this helps :)

OpenStudy (unklerhaukus):

A triangular pyramid has 4 faces. So, simply multiply the surface area of one face, by 4.

OpenStudy (anonymous):

Those were examples I gave you. So now for your question...120 x 4= 480. The triangle has 3 faces, plus the base so, 3+1=4. 4 times the surface area of the trianles faces would be....120 x 4 which equals 480.